# Stochastic Control Methods for Optimization

### v1

@@ 3.2 Regularization error
$\mathcal{H}(\mu^{\delta}|\gamma_{x_{0}})=\int_{\mathbb{R}^{d}}\ln\left(\frac{\mu^{\delta}(dx)}{\gamma_{x_{0}}(dx)}\right)\mu^{\delta}(dx)=-h(\mu^{\delta})-\int_{\mathbb{R}^{d}}\ln\left(\frac{\gamma_{x_{0}}(dx)}{dx}\right)\mu^{\delta}(dx),$
where $h(\mu^{\delta})=\int\ln\left(\frac{\mu^{\delta}(dx)}{dx}\right)\mu^{\delta}(dx)$ is the differential entropy. On the one hand, as conditioning reduces entropy, we have $h(\mu^{\delta})\geq h(\gamma_{0}^{\delta})$, where $\gamma_{0}^{\delta}=\mathcal{N}(0,\delta I_{d})$ is the Gaussian with mean 0 and covariance $\delta I_{d}$, which yields the upper bound
$h(\mu^{\delta})\leq-h(\gamma_{0}^{\delta})=-\frac{d}{2}\ln(2\pi e\delta).$
On the other hand, we estimate the cross-term:
$\int_{\mathbb{R}^{d}}\ln\left(\frac{\gamma_{x_{0}}(dx)}{dx}\right)\mu^{\delta}(dx)=\int_{\mathbb{R}^{d}}\left(\frac{|x-x_{0}|^{2}}{2}+\frac{d}{2}\ln(2\pi)\right)\mu^{\delta}(dx)\leq M_{2}(x_{0})+d\delta+\frac{d}{2}\ln(2\pi),$
where $M_{2}(x_{0})=\int|x-x_{0}|^{2}d\mu^{*}(x)$ is nothing but $\mathcal{W}_{2}(\delta_{x_{0}},\mu^{*})^{2}$.
@@
$\mathcal{H}(\mu^{\delta}|\gamma_{x_{0}})\leq C_{1}+\frac{d}{2}\ln\left(\frac{1}{\delta}\right)$ (3.16)
for some constant $C_{1}$ depending on $d$ and $\mathcal{W}_{2}(\delta_{x_{0}},\mu^{*})^{2}$.
Step 3. Substituting the estimates back into the upper bound for $\mathcal{V}_{\varepsilon}$ gives
$\mathcal{V}_{\varepsilon}-V_{0}\leq L_{G}\sqrt{d\delta}+\varepsilon\left(C_{1}+\frac{d}{2}\ln\left(\frac{1}{\delta}\right)\right).$
Setting $\delta=\varepsilon^{2}$ yields that
$\mathcal{V}_{\varepsilon}-V_{0}\leq C\varepsilon\ln\frac{1}{\varepsilon},$
where the constant $C$ depends on $d$, $L_{G}$, and $\mathcal{W}_{2}(\delta_{x_{0}},\mu^{*})$. ∎
@@ 3.3 Numerical approximation with examples
Here, we focus on the numerical approximation of the $N$-particle control problem (1.21) as an approximation to the regularized MFC problem. Observe that the $N$-particle control problem (1.21) constitutes a finite-dimensional stochastic control problem on $\mathbb{R}^{dN}$, in the same spirit as the setting studied in Section 2. Consequently, numerical procedures can be developed analogously to those in Section 2.3, by employing time discretization of the controlled SDEs with Monte Carlo simulations.
@@
    for each particle $i$, the empirical measure $\frac{1}{N}\sum_{j\neq i}^{N}\delta_{X^{j}_{t_{k}}}$ is used in the computation of $\beta^{i}$, reflecting the mean-field dependence;
@@
    although we may iterate the above scheme multiple times, each time we initialize the particles only through the terminal state of the previous iteration (without involving the mean)), basically because we are not searching for the mean but for the distribution.
@@
Algorithm 2 Stochastic Control for Optimization on $\mathcal{P}_{2}(\mathbb{R}^{d})$
@@
6:   Project Context: $\tilde{X}^{(k)}\leftarrow X_{t}^{(k)}+\sqrt{\tau}\zeta^{(k)}$ for $\zeta\sim\mathcal{N}(0,I)$; form $\tilde{\mu}_{T}\approx\frac{1}{N}\sum\delta_{\tilde{X}^{(k)}}$
@@
8:     Sample $Y_{i,j}\leftarrow X_{t}^{(i)}+\sqrt{\tau}\xi_{i,j}$ for $j=1..S$, $\xi\sim\mathcal{N}(0,I)$
@@
10:     Compute drift $\hat{\theta}_{i}\leftarrow\frac{1}{\tau}\left(\frac{\sum w_{i,j}Y_{i,j}}{\sum w_{i,j}}-X_{t}^{(i)}\right)$ (if $\tau>\delta$)
11:     Update $X_{t+\Delta t}^{(i)}\leftarrow X_{t}^{(i)}+\hat{\theta}_{i}\Delta t+\sqrt{\Delta t}Z^{(i)}$
@@
Then for each particle $i$, the independence of noises gives us the following simplified computation for $\beta^{i}$:
@@
We consider the minimization of the energy functional associated with Newtonian interaction in 2D:
$G(\mu)=\frac{1}{2}\iint_{\mathbb{R}^{2d}}\left(\frac{|x-y|^{2}}{2}-\ln(|x-y|)\right)\,d\mu(x)\,d\mu(y).$ (3.22)
@@
Figure 4: Convergence Rate of Energy Level. The plot of energy $G(\mu_{\varepsilon})$ versus $\frac{1}{N}$ shows a linear trend, confirming the $O(1/N)$ convergence rate predicted by Theorem 3.3.
@@
Appendix A Proof of Theorem 3.3
@@
The proof relies on the method of projection and the comparison principle for HJB equations. We construct an approximate solution to the finite-dimensional HJB equation by projecting the solution of the (infinite-dimensional) master equation onto the empirical manifold.
Step 1 Define the projected function $w^{N}_{\varepsilon}:[0,1]\times(\mathbb{R}^{d})^{N}\to\mathbb{R}$ as the value of the mean-field problem scaled by $N$, evaluated at the empirical measure:
$w^{N}_{\varepsilon}(t,\mathbf{x}):=N\cdot\mathcal{V}_{\varepsilon}\left(t,\frac{1}{N}\sum_{i=1}^{N}\delta_{x_{i}}\right),\quad\text{for }(t,\mathbf{x})\in[0,1]\times(\mathbb{R}^{d})^{N}.$ (A.1)
We calculate the derivatives of $w^{N}_{\varepsilon}$ with respect to the particle positions $x_{i}$ using the chain rule for Lions derivatives. For any $i\in\{1,\dots,N\}$:
$\partial_{x_{i}}w^{N}_{\varepsilon}(t,\mathbf{x})=N\cdot\frac{1}{N}(\partial_{\mu}\mathcal{V}_{\varepsilon})(t,\mu^{N}_{\mathbf{x}},x_{i})=\partial_{\mu}\mathcal{V}_{\varepsilon}(t,\mu^{N}_{\mathbf{x}},x_{i}).$ (A.2)
For the second derivatives (Hessian), we have two terms coming from the dependence on the measure:
$\Delta_{x_{i}}w^{N}_{\varepsilon}(t,\mathbf{x})=\text{div}_{x}(\partial_{\mu}\mathcal{V}_{\varepsilon})(t,\mu^{N}_{\mathbf{x}},x_{i})+\frac{1}{N}\text{Tr}\left(\partial^{2}_{\mu\mu}\mathcal{V}_{\varepsilon}(t,\mu^{N}_{\mathbf{x}},x_{i},x_{i})\right),$ (A.3)
@@
$\mathcal{L}^{N}[\phi]=-\partial_{t}\phi-\frac{1}{2}\sum_{i=1}^{N}\Delta_{x_{i}}\phi+\frac{1}{2\varepsilon}\sum_{i=1}^{N}|\partial_{x_{i}}\phi|^{2}.$
@@
$\mathcal{L}^{N}[w^{N}_{\varepsilon}]$ $=-N\partial_{t}\mathcal{V}_{\varepsilon}-\frac{1}{2}\sum_{i=1}^{N}\left(\text{div}_{y}(\partial_{\mu}\mathcal{V}_{\varepsilon})(x_{i})+\frac{1}{N}\text{Tr}(\partial^{2}_{\mu\mu}\mathcal{V}_{\varepsilon})(x_{i},x_{i})\right)$
$\quad+\frac{1}{2\varepsilon}\sum_{i=1}^{N}|\partial_{\mu}\mathcal{V}_{\varepsilon}(x_{i})|^{2}.$

### v2

@@ 3.2 Regularization error
$\mathcal{H}(\mu^{\lambda}|\gamma_{x_{0}})=\int_{\mathbb{R}^{d}}\ln\left(\frac{\mu^{\lambda}(dx)}{\gamma_{x_{0}}(dx)}\right)\mu^{\lambda}(dx)=-h(\mu^{\lambda})-\int_{\mathbb{R}^{d}}\ln\left(\frac{\gamma_{x_{0}}(dx)}{dx}\right)\mu^{\lambda}(dx),$
where $h(\mu^{\lambda})=-\int_{\mathbb{R}^{d}}\ln\left(\frac{\mu^{\lambda}(dx)}{dx}\right)\mu^{\lambda}(dx)$ is the differential entropy. On the one hand, as conditioning reduces entropy, we have $h(\mu^{\lambda})\geq h(\gamma_{0}^{\lambda})$, where $\gamma_{0}^{\lambda}=\mathcal{N}(0,\lambda I_{d})$ is the Gaussian with mean 0 and covariance $\lambda I_{d}$, which yields the upper bound
$h(\mu^{\lambda})\leq-h(\gamma_{0}^{\lambda})=-\frac{d}{2}\ln(2\pi e\lambda).$
On the other hand, we estimate the cross-term
$\int_{\mathbb{R}^{d}}\ln\left(\frac{\gamma_{x_{0}}(dx)}{dx}\right)\mu^{\lambda}(dx)=\int_{\mathbb{R}^{d}}\left(\frac{|x-x_{0}|^{2}}{2}+\frac{d}{2}\ln(2\pi)\right)\mu^{\lambda}(dx)\leq M_{2}(x_{0})+d\lambda+\frac{d}{2}\ln(2\pi),$
where $M_{2}(x_{0})=\int|x-x_{0}|^{2}d\mu^{*}(x)$ is nothing but $|\mathcal{W}_{2}(\delta_{x_{0}},\mu^{*})|^{2}$.
@@
$\mathcal{H}(\mu^{\lambda}|\gamma_{x_{0}})\leq C_{1}+\frac{d}{2}\ln\left(\frac{1}{\lambda}\right)$ (3.9)
for some constant $C_{1}$ depending on $d$ and $\mathcal{W}_{2}(\delta_{x_{0}},\mu^{*})$.
Step 3. Substituting the estimates back into the upper bound for $\mathcal{V}_{\varepsilon}(0,\delta_{x_{0}})$ gives
$\mathcal{V}_{\varepsilon}(0,\delta_{x_{0}})-\mathcal{V}_{0}\leq L_{G}(d\lambda)^{\alpha/2}+\varepsilon\left(C_{1}+\frac{d}{2}\ln\left(\frac{1}{\lambda}\right)\right).$
Setting $\lambda=\frac{\varepsilon^{2/\alpha}}{d}$ yields that
$\mathcal{V}_{\varepsilon}(0,\delta_{x_{0}})-\mathcal{V}_{0}\leq C\varepsilon\ln\left(\frac{1}{\varepsilon}\right),$
where the constant $C$ depends on $d$, $L_{G}$, $\alpha$, and $\mathcal{W}_{2}(\delta_{x_{0}},\mu^{*})$. ∎
@@ 3.3 Numerical approximation with examples
Here, we focus on the numerical approximation of the $N$-particle control problem (1.21) as an approximation to the regularized MFC problem. Observe that the $N$-particle control problem (1.21) constitutes a finite-dimensional stochastic control problem on $\mathbb{R}^{dN}$, in the same spirit as the setting studied in Section 2. Consequently, numerical approximations can be developed analogously to those in Section 2.3, by employing time discretization of the controlled SDEs with Monte Carlo simulations.
@@
    for each particle $i$, the empirical measure $\frac{1}{N}\sum_{j=1}^{N}\delta_{X^{j}_{t_{k}}}$ is used in the computation of $\beta^{i}$, reflecting the mean-field dependence;
@@
    while the procedure may be repeated over multiple outer iterations, the re-initialization at each iteration is taken directly from the terminal particle configuration of the preceding iteration (without an additional averaging/coupling step), since the objective is to approximate an optimizing distribution rather than a single point estimator.
@@
Algorithm 2 Stochastic Control Method for Optimization on $\mathcal{P}_{2}(\mathbb{R}^{d})$
@@
6:   Project Context: $\tilde{X}^{(k)}\leftarrow X_{t}^{(k)}+\sqrt{\tau}\zeta^{(k)}$ for $\zeta^{(k)}\sim\mathcal{N}(0,I)$; form $\tilde{\mu}_{T}\approx\frac{1}{N}\sum\delta_{\tilde{X}^{(k)}}$
@@
8:     Sample $Y_{i,j}\leftarrow X_{t}^{(i)}+\sqrt{\tau}\xi_{i,j}$, for $j=1..S$
@@
10:     Compute drift $\hat{\theta}_{i}\leftarrow\frac{1}{\tau}\left(\frac{\sum w_{i,j}Y_{i,j}}{\sum w_{i,j}}-X_{t}^{(i)}\right)$
11:     Update $X_{t+\Delta t}^{(i)}\leftarrow X_{t}^{(i)}+\hat{\theta}_{i}\Delta t+\sqrt{\Delta t}Z^{(i)}$, for $Z^{(i)}\sim\mathcal{N}(0,I)$
@@
Then for each particle $i$, the independence of noises gives us the following simplified computation
@@
Consider the minimization of the energy functional associated with Newtonian interaction in 2D:
$G(\mu)=\iint_{\mathbb{R}^{2d}}\left(\frac{|x-y|^{2}}{2}-\ln(|x-y|)\right)\,d\mu(x)\,d\mu(y).$ (3.15)
@@
Figure 4: Convergence Rate of Value Errors. The plot of value error versus $\frac{1}{N}$ shows a linear trend, confirming the $O(1/N)$ convergence rate predicted by Theorem 3.3.
@@
Appendix A Proof of Proposition 3.2
Proof.
Under the assumptions (a) (Regularity) and (b) (Convexity) in Proposition 3.2, the stochastic Pontryagin maximum principle with both necessary and sufficient conditions established in [11, Theorems 4.5, 4.6 & 5.1] implies that, for each $(t,\mu)\in[0,T)\times\mathcal{P}_{2}(\mathbb{R}^{d})$, $\xi\sim\mu$, the optimal control $\theta^{*}$ exists uniquely, and the optimal state process $X^{t,\xi;\theta^{*}}$ satisfies uniquely the following FBSDE:
$\begin{cases}dX^{t,\xi;\theta^{*}}_{s}=-\frac{1}{\varepsilon}Y^{t,\xi;\theta^{*}}_{s}\,ds+dW_{s},\quad X^{t,\xi;\theta^{*}}_{t}=\xi\sim\mu,\\ dY^{t,\xi;\theta^{*}}_{s}=Z^{t,\xi;\theta^{*}}_{s}\,dW_{s},\quad Y^{t,\xi;\theta^{*}}_{1}=\partial_{\mu}G(\mathcal{L}(X^{t,\xi;\theta^{*}}_{1}))(X^{t,\xi;\theta^{*}}_{1}),\\ \theta^{*}_{s}=-\frac{1}{\varepsilon}Y^{t,\xi;\theta^{*}}_{s},\quad\forall\,s\in[t,T],\end{cases}$ (A.1)
which, combined with the theories in [16, Section 5.3 and Theorem 5.7], implies that the value function $\mathcal{V}_{\varepsilon}$ is a classical solution to the master equation (1.20) satisfying Assumption 3.1 and that there holds the decoupling relation $Y^{t,\xi;\theta^{*}}_{s}=\partial_{\mu}\mathcal{V}_{\varepsilon}(s,\mathcal{L}(X^{t,\xi;\theta^{*}}_{s}))(X^{t,\xi;\theta^{*}}_{s})$ for all $s\in[t,T]$. By Lemma 3.1, the classical solution is unique.
It remains to justify the $\varepsilon$-uniform boundedness estimate (3.1). Inspired by [12, Chapter 2, Volume II], we may derive the higher-order regularity of $\mathcal{V}_{\varepsilon}$ by differentiating FBSDE (A.1).
Let $\xi\sim\mu$. First, differentiating the FBSDE (A.1) with respect to the spatial initial condition $x$ gives (cf.[12, Eq (5.65), Volume II])
$\begin{cases}d\big(\partial_{x}X^{t,x;\theta^{*}}_{s}\big)=-\dfrac{1}{\varepsilon}\,\partial_{x}Y^{t,x;\theta^{*}}_{s}\,ds,\qquad\partial_{x}X^{t,x;\theta^{*}}_{t}=I_{d},\\[6.00006pt] d\big(\partial_{x}Y^{t,x;\theta^{*}}_{s}\big)=\partial_{x}Z^{t,x;\theta^{*}}_{s}\,dW_{s},\qquad\partial_{x}Y^{t,x;\theta^{*}}_{1}=\partial_{x}\partial_{\mu}G\!\left(\mathcal{L}\!\left(X^{t,\xi;\theta^{*}}_{1}\right)\right)\!\left(X^{t,x,\xi;\theta^{*}}_{1}\right)\,\partial_{x}X^{t,x;\theta^{*}}_{1},\end{cases}$
where $X^{t,\xi;\theta^{*}}_{1}$ satisfies the original FBSDE (A.1) while $X_{s}^{t,x,\xi;\theta^{*}}$ together with the pair $(Y_{s}^{t,x,\xi;\theta^{*}},Z_{s}^{t,x,\xi;\theta^{*}})$ satisfies FBSDE
$\begin{cases}dX^{t,x,\xi;\theta^{*}}_{s}=-\frac{1}{\varepsilon}Y^{t,x,\xi;\theta^{*}}_{s}\,ds+dW_{s},\quad X^{t,x,\xi;\theta^{*}}_{t}=x,\\ dY^{t,x,\xi;\theta^{*}}_{s}=Z^{t,x,\xi;\theta^{*}}_{s}\,dW_{s},\quad Y^{t,x,\xi;\theta^{*}}_{1}=\partial_{\mu}G(\mathcal{L}(X^{t,\xi;\theta^{*}}_{1}))(X^{t,x,\xi;\theta^{*}}_{1}).\end{cases}$
Applying Itô’s formula to the scalar product $\langle\partial_{x}X^{t,x;\theta^{*}}_{s},\partial_{x}Y_{s}^{t,x;\theta^{*}}\rangle$ and taking conditional expectation, we obtain
$\langle\partial_{x}X^{t,x;\theta^{*}}_{s},\partial_{x}Y_{s}^{t,x;\theta^{*}}\rangle$ $=\mathbb{E}\Big[\underbrace{\langle\partial_{x}X_{1}^{t,x;\theta^{*}},\,\partial_{x}\partial_{\mu}G(\mathcal{L}(X^{t,\xi;\theta^{*}}_{1}))(X^{t,x,\xi;\theta^{*}}_{1})\partial_{x}X_{1}^{t,x;\theta^{*}}\rangle}_{\geq 0\text{ by Coercivity}}+\frac{1}{\varepsilon}\int_{s}^{1}|\partial_{x}Y_{r}^{t,x;\theta^{*}}|^{2}\,dr\big|\mathcal{F}_{s}\Big],$
$\geq 0,\text{ a.s. }\qquad\forall\,s\in[t,1].$ (A.2)
Applying the chain rule further yields
$d(|\partial_{x}X^{t,x;\theta^{*}}_{s}|^{2})=2\langle\partial_{x}X_{s}^{t,x;\theta^{*}},d(\partial_{x}X_{s}^{t,x;\theta^{*}})\rangle=-\frac{2}{\varepsilon}\langle\partial_{x}X_{s}^{t,x;\theta^{*}},\partial_{x}Y_{s}^{t,x;\theta^{*}}\rangle\,ds\leq 0.$ (A.3)
Thus, it holds that $|\partial_{x}X_{s}^{t,x;\theta^{*}}|\leq|I_{d}|=\sqrt{d}$ for all $s\in[t,1]$. Looking at $\partial_{x}Y^{t,x;\theta^{*}}$, we have
$|\partial_{x}\partial_{\mu}\mathcal{V}_{\varepsilon}(t,\mu)(x)|=|\partial_{x}Y_{t}^{t,x;\theta^{*}}|$ $=\left|\mathbb{E}\left[\partial_{x}(\partial_{\mu}G(\mathcal{L}(X^{t,\xi;\theta^{*}}_{1}))(X^{t,\xi;\theta^{*}}_{1}))\cdot\partial_{x}X^{t,x;\theta^{*}}_{1}\big|\mathcal{F}_{t}\right]\right|$
$\leq K_{G}\mathbb{E}[|\partial_{x}X^{t,x;\theta^{*}}_{1}||\mathcal{F}_{t}]$
$\leq K_{G}\sqrt{d}.$
Therefore, we obtain the desired bound for $\partial_{x}\partial_{\mu}\mathcal{V}_{\varepsilon}$.
Let us differentiate the FBSDE (A.1) with respect to the measure $\mu$. Let $\phi\in L^{2}(\mathbb{R}^{d},\mu;\mathbb{R}^{d})$ be a vector field representing the direction of the perturbation. This induces a perturbation of the initial random variable $\xi$ given by $\delta X^{\phi}_{t}=\phi(\xi)\in L^{2}(\Omega;\mathbb{R}^{d})$. The tangent process $(\delta X_{s},\delta Y_{s},\delta Z_{s})$ is defined as the $L^{2}(\Omega)$-limit of the difference quotients:
where $\theta^{r}$ is the optimal control associated to the initial condition $\xi+r\phi(\xi)$. Differentiating the FBSDE coefficients leads to the following linearized system
$\begin{cases}d(\delta X_{s})=-\frac{1}{\varepsilon}\delta Y_{s}\,ds,\quad\delta X_{t}=\phi(\xi),\\ d(\delta Y_{s})=\delta Z_{s}\,dW_{s},\\ \delta Y_{1}=\partial_{x}\partial_{\mu}G(\mathcal{L}(X^{t,\xi;\theta^{*}}_{1}))(X^{t,\xi;\theta^{*}}_{1})\cdot\delta X_{1}+\tilde{\mathbb{E}}[\partial^{2}_{\mu\mu}G(\mathcal{L}(X^{t,\xi;\theta^{*}}_{1}))(X^{t,\xi;\theta^{*}}_{1})(\tilde{X}^{t,\xi;\theta^{*}}_{1})\cdot\delta\tilde{X}_{1}].\end{cases}$ (A.4)
In a similar manner to (A.2)-(A.3), applying Itô’s formulas and the convexity assumption, we obtain
$|\delta X_{1}|\leq|\phi(\xi)|\text{ a.s., }$ (A.5)
and further,
$\mathbb{E}[|\delta Y_{t}|]$ $=\mathbb{E}\left[\left|\mathbb{E}\left[\partial_{x}\partial_{\mu}G(\mathcal{L}(X^{t,\xi;\theta^{*}}_{1}))(X^{t,\xi;\theta^{*}}_{1})\cdot\delta X_{1}+\tilde{\mathbb{E}}[\partial^{2}_{\mu\mu}G(\mathcal{L}(X^{t,\xi;\theta^{*}}_{1}))(X^{t,\xi;\theta^{*}}_{1})(\tilde{X}^{t,\xi;\theta^{*}}_{1})\cdot\delta\tilde{X}_{1}]\big|\mathcal{F}_{t}\right]\right|\right]$
$\leq K_{G}\mathbb{E}[|\delta X_{1}|]+K_{G}\tilde{\mathbb{E}}[|\delta\tilde{X}_{1}|]$
$\leq 2K_{G}\mathbb{E}[|\phi(\xi)|].$ (A.6)
Then, we investigate the directional derivative of $\mathcal{V}_{\varepsilon}$ in the direction $\phi$. In view of [12, Eq. (5.59), Volume II], denote the associated directional derivatives by $\delta_{\phi}X_{s}$, $\delta_{\phi}Y_{s}$, $\delta_{\phi}Z_{s}$, which satisfy the following FBSDE
$\begin{cases}d(\delta_{\phi}X_{s})=-\frac{1}{\varepsilon}\delta_{\phi}Y_{s}\,ds,\qquad\delta_{\phi}X_{t}=0,\\ d(\delta_{\phi}Y_{s})=\delta_{\phi}Z_{s}\,dW_{s},\\ \delta_{\phi}Y_{1}=\partial_{x}\partial_{\mu}G(\mathcal{L}(X^{t,\xi;\theta^{*}}_{1}))(X^{t,x,\xi;\theta^{*}}_{1})\cdot\delta_{\phi}X_{1}+\tilde{\mathbb{E}}[\partial^{2}_{\mu\mu}G(\mathcal{L}(X^{t,\xi;\theta^{*}}_{1}))(X^{t,x,\xi;\theta^{*}}_{1})(\tilde{X}^{t,\xi;\theta^{*}}_{1})\cdot\delta\tilde{X}_{1}].\end{cases}$ (A.7)
Applying Itô’s formula to $\langle\delta_{\phi}X_{s},\delta_{\phi}Y_{s}\rangle$ and taking expectations give
$\mathbb{E}\left[\langle\delta_{\phi}X_{1},\delta_{\phi}Y_{1}\rangle\right]$ $=-\frac{1}{\varepsilon}\mathbb{E}\left[\int_{t}^{1}|\delta_{\phi}Y_{s}|^{2}ds\right]+\mathbb{E}\left[\langle\delta_{\phi}Y_{t},\,\delta_{\phi}X_{t}\rangle\right]=-\frac{1}{\varepsilon}\mathbb{E}\left[\int_{t}^{1}|\delta_{\phi}Y_{s}|^{2}ds\right]\leq 0.$
On the other hand, the boundedness of $\partial_{\mu\mu}^{2}G$ and the coercivity assumption yield that
$0\geq\mathbb{E}\left[\langle\delta_{\phi}X_{1},\delta_{\phi}Y_{1}\rangle\right]$ $=\mathbb{E}\left[\langle\delta_{\phi}X_{1},\partial_{x}\partial_{\mu}G(\mathcal{L}(X^{t,\xi;\theta^{*}}_{1}))(X^{t,x,\xi;\theta^{*}}_{1})\cdot\delta_{\phi}X_{1}\rangle\right]$
$\qquad+\mathbb{E}\left[\langle\delta_{\phi}X_{1},\tilde{\mathbb{E}}[\partial^{2}_{\mu\mu}G(\mathcal{L}(X^{t,\xi;\theta^{*}}_{1}))(X^{t,x,\xi;\theta^{*}}_{1})(\tilde{X}^{t,\xi;\theta^{*}}_{1})\cdot\delta\tilde{X}_{1}]\rangle\right]$
$\geq\kappa\,\mathbb{E}\left[|\delta_{\phi}X_{1}|^{2}\right]-K_{G}\,\mathbb{E}\left[|\delta_{\phi}X_{1}|\right]\cdot\tilde{\mathbb{E}}\left[|\delta\tilde{X}_{1}|\right],$
which together with (A.5) implies that $\sqrt{\mathbb{E}\left[|\delta_{\phi}X_{1}|^{2}\right]}\leq\frac{K_{G}}{\kappa}\mathbb{E}\left[|\phi(\xi)|\right]$. Hence, by the BSDE for $\delta_{\phi}Y_{t}$, we have
$|\mathbb{E}\left[\partial^{2}_{\mu\mu}\mathcal{V}_{\varepsilon}(t,\mu)(x)(\xi)\cdot\phi(\xi)\right]|=|\mathbb{E}[\delta_{\phi}Y_{t}]|$
$=\left|\mathbb{E}\left[\partial_{x}\partial_{\mu}G(\mathcal{L}(X^{t,\xi;\theta^{*}}_{1}))(X^{t,x,\xi;\theta^{*}}_{1})\cdot\delta_{\phi}X_{1}+\tilde{\mathbb{E}}[\partial^{2}_{\mu\mu}G(\mathcal{L}(X^{t,\xi;\theta^{*}}_{1}))(X^{t,x,\xi;\theta^{*}}_{1})(\tilde{X}^{t,\xi;\theta^{*}}_{1})\cdot\delta\tilde{X}_{1}]\big|\mathcal{F}_{t}\right]\right|$
$\leq K_{G}\mathbb{E}[|\delta_{\phi}X_{1}|]+K_{G}\tilde{\mathbb{E}}[|\delta\tilde{X}_{1}|]$
$\leq\left(\frac{(K_{G})^{2}}{\kappa}+K_{G}\right)\mathbb{E}[|\phi(\xi)|].$
In view of the arbitrariness of $\phi$, we obtain the desired bound for $\partial^{2}_{\mu\mu}\mathcal{V}_{\varepsilon}$. ∎
Appendix B Proof of Theorem 3.3
@@
The proof relies on the method of projection and the comparison principle for parabolic equations.
Step 1 Define the projected function $w^{N}_{\varepsilon}:[0,1]\times(\mathbb{R}^{d})^{N}\to\mathbb{R}$ as the value of the mean-field problem scaled by $N$, evaluated at the empirical measure $\mu^{N}_{\mathbf{x}}=\frac{1}{N}\sum_{i=1}^{N}\delta_{x_{i}}$, i.e.,
$w^{N}_{\varepsilon}(t,\mathbf{x}):=N\cdot\mathcal{V}_{\varepsilon}\left(t,\frac{1}{N}\sum_{i=1}^{N}\delta_{x_{i}}\right),\quad\text{for }(t,\mathbf{x})\in[0,1]\times(\mathbb{R}^{d})^{N}.$ (B.1)
We calculate the derivatives of $w^{N}_{\varepsilon}$ with respect to the particle positions $x_{i}$ using the chain rule for Lions derivatives. For each $i\in\{1,\dots,N\}$,
$\partial_{x_{i}}w^{N}_{\varepsilon}(t,\mathbf{x})=N\cdot\frac{1}{N}(\partial_{\mu}\mathcal{V}_{\varepsilon})(t,\mu^{N}_{\mathbf{x}})(x_{i})=\partial_{\mu}\mathcal{V}_{\varepsilon}(t,\mu^{N}_{\mathbf{x}})(x_{i}).$ (B.2)
For the second derivatives (Hessian), we have
$\Delta_{x_{i}}w^{N}_{\varepsilon}(t,\mathbf{x})=\text{div}_{x}(\partial_{\mu}\mathcal{V}_{\varepsilon})(t,\mu^{N}_{\mathbf{x}})(x_{i})+\frac{1}{N}\text{Tr}\left(\partial^{2}_{\mu\mu}\mathcal{V}_{\varepsilon}(t,\mu^{N}_{\mathbf{x}})(x_{i})(x_{i})\right),$ (B.3)
@@
$\mathcal{L}^{N}[\phi](t,\mathbf{x})=-\partial_{t}\phi(t,\mathbf{x})-\frac{1}{2}\sum_{i=1}^{N}\Delta_{x_{i}}\phi(t,\mathbf{x})+\frac{1}{2\varepsilon}\sum_{i=1}^{N}|\partial_{x_{i}}\phi(t,\mathbf{x})|^{2}.$
@@
$\mathcal{L}^{N}[w^{N}_{\varepsilon}](t,\mathbf{x})$ $=-N\partial_{t}\mathcal{V}_{\varepsilon}(t,\mu^{N}_{\mathbf{x}})-\frac{1}{2}\sum_{i=1}^{N}\left(\text{div}_{x_{i}}(\partial_{\mu}\mathcal{V}_{\varepsilon})(t,\mu^{N}_{\mathbf{x}})(x_{i})+\frac{1}{N}\text{Tr}(\partial^{2}_{\mu\mu}\mathcal{V}_{\varepsilon})(t,\mu^{N}_{\mathbf{x}})(x_{i})(x_{i})\right)$
$\quad+\frac{1}{2\varepsilon}\sum_{i=1}^{N}|\partial_{\mu}\mathcal{V}_{\varepsilon}(t,\mu^{N}_{\mathbf{x}})(x_{i})|^{2}.$